OpenStudy (anonymous):

Which of the following is the solution to the equation (sqrt5)^8n = 125^(n + 4)

OpenStudy (anonymous):

\[\sqrt{5}^{8n}=5^{4n}\] \[125^{n+4}=5^{3(n+4)}\]

OpenStudy (anonymous):

so your job is to solve \[4n=3(n+4)\]

OpenStudy (anonymous):

i get \[n=12\]